OpenStudy (milo123):

MEDAL What is the length, in units, of the hypotenuse of a right triangle if each of the two legs is 3 units? Square root of 18 units Square root of 27 units 3 units 6 units

OpenStudy (milo123):

3^2 x 3^2 = 81

OpenStudy (milo123):

@mathmate @Mehek14

OpenStudy (jango_in_dtown):

By Pythagorus theorem , the answer is sqrt( 3^2+3^2)= sqrt 18

OpenStudy (milo123):

oops i accidentally did multiplication

OpenStudy (jango_in_dtown):

its ok. :)

OpenStudy (andrewishere):

Ok so lets draw this

OpenStudy (andrewishere):

So use the pythagorean formula \[a^2 +b^2 + c^2\] for the following 

OpenStudy (andrewishere):

a = 3 b = 3 c = ?

OpenStudy (andrewishere):

Plug into the pythagorean Theorem: \[(3)^2+(3)^2 = c^2\]

OpenStudy (andrewishere):

equals

OpenStudy (andrewishere):

\[9+9 = c^2\]

OpenStudy (andrewishere):

equals

OpenStudy (andrewishere):

\[18 = c^2\]

OpenStudy (andrewishere):

To isolate c, find the square root of both sides

OpenStudy (andrewishere):

\[\sqrt{18}= \sqrt{c^2}\]

OpenStudy (andrewishere):

So...

OpenStudy (andrewishere):

\[c = \sqrt{18}\] or simplified \[c = 3\sqrt{2}\]

OpenStudy (milo123):

thanks

OpenStudy (andrewishere):

Sorry it was so fast, but I am in school and I have to go to my next class. I hoped this helped. Any more questions, just message me and I will get to it ASAP